what size container do you need to hold 4.97 mol of N2 gas at STP


1
Expert's answer
2021-01-22T06:23:43-0500

Solution:

1st Solution Path:

At STP, one mole of any gas occupies a volume of 22.4 dm3.

Thus 4.97 mol of N2 occupies:

(4.97 mol × 22.4 dm3) / 1 mol = 111.33 dm3 = 111.3 dm3

V(container) = 111.3 dm3


2nd Solution Path:

STP is defined as a temperature of 273.15 K and an absolute pressure of 101.325 kPa.


The Ideal Gas equation (PV = nRT) can be used.


P = 101.325 kPa

V = unknown

n = 4.97 mol

R = 8.314 J⋅K⋅mol−1

T = 273.15 K


Hence,

V = nRT / P

V = (4.97 mol × 8.314 J⋅K⋅mol−1 × 273.15 K) / (101.325 kPa) = 111.39 dm3 = 111.4 dm3

V (container) = 111.4 dm3


Answer: The size container is 111.3 dm3.